A car accelerates uniformly, from rest, at 2 m.s for 20s. it then travels at constant velocity for 60 s and finally brakes uniformly to rest in 8 s. Calculate the total distance covered.


1
Expert's answer
2021-02-22T10:26:41-0500
"v=at_1=(2)(20)=40\frac{m}{s}\\s=0.5(2)(20)^2+(40)(60)+0.5(8)(40)=2960\ m"